# Neural-NPV Control: Learning Parameter-Dependent Controllers and Lyapunov Functions with Neural Networks

### 17 Mar 2026 v1

@@ I Introduction
A nonlinear parameter-varying (NPV) system is a nonlinear system whose behavior explicitly depends on external time-varying parameters. This concept extends the linear parameter-varying (LPV) framework [21, 19], by allowing for the dynamics to have a nonlinear instead of a linear structure. Many real-world nonlinear time-varying systems can be effectively represented as NPV systems. Examples include rockets with decreasing mass due to fuel consumption, ground vehicles subjected to variable friction coefficients, and aircraft experiencing disturbances that change over time. Similar to the LPV framework [21, 19, 14], representing a nonlinear time-varying system as an NPV system facilitates the design of gain-scheduled nonlinear controllers that adjust the control strategy dynamically based on real-time measurements and estimates of the scheduling parameters.
Recent studies have explored the synthesis of nonlinear parameter-dependent (PD) controllers for NPV systems using SOS programming [10, 12, 30]. However, SOS programming is not scalable to high-dimensional systems. Additionally, SOS-based nonlinear synthesis often imposes a restrictive structure on the Lyapunov functions to avoid a non-convex problem, leading to potentially conservative results. To overcome the limitations of SOS-based NPV synthesis, this paper investigates the use of NNs to learn stabilizing control laws for NPV systems and Lyapunov functions for certifying the stability of the closed-loop system.
@@ I-A Related work
SOS and its use in NPV synthesis: One of the most widely used methods for nonlinear analysis and synthesis is SOS optimization [15]. It has been applied to jointly synthesize nonlinear parameter-dependent (PD) controllers and PD Lyapunov functions for NPV systems [10, 12, 30]. However, SOS optimization-based methods suffer from two major limitations: (i) Restricted scalability: The computational complexity grows rapidly with the dimensionality of the system and the degree of the polynomial basis used, rendering them impractical for high-dimensional systems [2]. (ii) Conservative performance due to structural restrictions: SOS-based nonlinear synthesis typically imposes structures on certificate functions to avoid a nonconvex problem, leading to conservative performance. For example, SOS-based Lyapunov synthesis typically requires that the Lyapunov functions do not depend on states whose derivatives are directly affected by control inputs [16, 13, 28, 20, 30].
@@
Learning certified control: There has been an increasing trend for learning a control certificate (alongside a controller) using NNs (NNs) [7]. Such a neural certificate can be a Lyapunov function for certifying the stability of an equilibrium point [17, 3, 5, 6], a contraction metric for certifying stability about trajectories [22, 24, 23], or a barrier function for certifying set invariance [18, 9], among others. Compared to traditional methods for nonlinear synthesis, e.g., based on sum-of-squares (SOS) programming [15], certified neural control can be applied to general nonlinear systems (in contrast to polynomial systems for SOS programming-based methods), and is scalable to high-dimensional systems. Nevertheless, existing work typically consider time-invariant dynamics [8]. To the best of our knowledge, there is no study on learning-based control for NPV systems.
@@ II Problem settings and preliminaries
$\theta(t)\in\Theta\;\;\text{and}\;\;\dot{\theta}(t)\in{\Omega},\quad\forall\,t\geq 0$ (3)
@@
[11, Definition 4.5] The equilibrium point $x^{*}=0$ of the closed-loop system 7 is exponentially stable if there exists positive constaints c,kc,~k, and $\lambda$ such that
@@
According to Lyapunov stability theory (as demonstrated in the proof of Theorem 1 in [30]), if there exists a PD controller $u=\pi(x,\theta)$ and a PD continuously differentiable function $V(x,\theta)\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}$ such that for all $x\in\mathcal{X}$ with $\mathcal{X}$ being a domain of interest,
@@
$\dot{V}(x,{\theta},\dot{\theta})\!=\!\frac{\partial V}{\partial x}\cdot\bar{f}+\frac{\partial V}{\partial\theta}\cdot\dot{\theta}=\frac{\partial V}{\partial x}\cdot f(x,\pi(x,\theta))+\frac{\partial V}{\partial\theta}\cdot\dot{\theta},$ (10)
@@ III Neural-NPV: Learning PD Controllers and Lyapunov functions
III-A Parameterizing the PD Lyapunov function and PD controller
@@
$\tilde{u}(x,\theta)=\pi_{\textup{NN}}(x,\theta)-\pi_{\textup{NN}}(x^{*},\theta)+\tilde{u}(x^{*},\theta)$ (14)
To enforce the constraint 2, we use tanh in the output layer of the controller $\pi_{\textup{NN}}$ to restrict the normalized output within [1,1][-1,1]. $\tilde{u}(x,\theta)$ and $\tilde{u}(x^{*},\theta)$ are normalized value of $u(x,\theta)$ and $u(x^{*},\theta)$, respectively. Rather than using projection or clipping, which introduces complications in gradient-based training, we employ a smooth, differentiable activation at the output, which empirically improves the training process. We scale the normalized output to actual physical control input using an affine de-normalization to get the final controller input:
@@ III-B The Neural-NPV algorithm
Within Neural-NPV, we jointly train the Lyapunov network $\phi_{\textup{NN}}$ and controller network $\pi_{\textup{NN}}$ using a two stage procedure. In Stage I, we use counterexample guided synthesis to obtain a working controller and a Lyapunov function candidate over the full domain $\mathcal{X}\times\Theta$, where $\mathcal{X}$ is the region of interest. In Stage II, we refine the controller and Lyapunov networks using a level-set-guided loss to certify a robust ROA and a surrogate loss that promotes the growth of the robust ROA. The algorithm is summarized in 35.
@@ III-B1 Stage I: Counterexample-guided joint synthesis
Existing work on learning certified control typically first generates samples from the state space, use these samples to evaluate the violation of the Lyapunov functions and optimizes the network parameters to minimize the violation. We follow this idea. However, here, we need to sample from the space of $x$, and $\theta$, and $\dot{\theta}$ i.e., $\mathcal{X}$, $\Theta$, and $\Omega$. At each iteration, we generate $N$ number of $x$-$\theta$ pairs, $(x^{1},\theta^{1}),\ (x^{N},\theta^{N})$ where $x^{i}\in\mathcal{X}$ and $\theta^{i}\in\Theta$ for $i=1,\dots,N$. We then use CEGIS (Counter-example guided Inductive Synthesis) [1] framework to iteratively generate adversarial samples $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and train the networks. Following [25, 27], we use the projected gradient descent (PGD) method to generate the worst adversarial examples that violate the following objective:
$\hskip-11.38109ptL_{\dot{V}}=\sum_{\dot{\theta}\in\textup{Ver}(\Omega)}\!\dot{V}(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}},\dot{\theta})+k_{3}\|x^{i}_{\textup{adv}}\|^{2},$ (16)
By maximizing 16, we aim to find adversarial samples that have the maximum violation of the condition 9c. Since 9a and 9b has already been satisfied by construction as explained in Section III-A, we only need to evaluate 9c. At each $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$, we evaluate the $V(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and $u(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and thus assess whether the Lyapunov conditions in 9 are violated at $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and if so, the level of violation. These adversarial examples are added to a buffer and then used to optimize the following empirical Lyapunov loss :
$\hskip-11.38109ptL_{lya}\!=\!\frac{1}{N}\sum_{i=1}^{N}\sum_{\dot{\theta}\in\textup{Ver}(\Omega)}\!\!\max\left(0,\dot{V}(x^{i}_{\textup{adv}},\theta^{i}_{adv},\dot{\theta})\!+\!k_{3}\|x^{i}_{\textup{adv}}\|^{2}\right),$ (17)
@@ III-B2 Stage II: Level set-guided ROA refinement
$\Lambda^{\rho}_{V}=\left\{x\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\max_{\theta\in\Theta}\,V(x,\theta)\leq\rho\right\}.$ (18)
@@
$\hskip-5.69054ptL_{\dot{V};\rho}\!=\frac{1}{N}\sum_{i=1}^{N}\!\min\left(L_{\dot{V}}(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}},\dot{\theta}),\rho\!-\!\max_{\theta\in\Theta}V(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})\right)$ (19)
@@
$L_{\Lambda^{\rho}_{V}}=\frac{1}{N}\sum_{i=1}^{N}\max_{\theta\in\Theta}V(x^{i}_{\textup{{candidate}}},\theta^{i}_{\textup{{candidate}}})-\rho$ (20)
Equation 20 indirectly promotes the growth of the level-set $\Lambda^{\rho}_{V}$. We sample uniformly from the state space to get ($x_{\textup{candidate}},\theta_{\textup{candidate}}$) that are used to compute 20. This surrogate loss ensures the robust level-set $\Lambda^{\rho}_{V}$ does not shrink to a neighborhood of the origin, and cover the region of interest as much as possible. The cost function of this refinement stage consists of both these objective functions. These two objectives plays a complementary role, 19 certifies $\Lambda^{\rho}_{V}$ as a valid robust ROA while 20 trying to maximize it’s volume.
III-B3 Initialization of the control network
Before entring Stage I, we initialize our PD controller to enhance the training process. As mentioned in the neural control literature [3], finding a feasible neural controller from random initialization of network parameters is difficult, as the stochastic gradient descent used in training is inherently local. Existing work often used LQR gains computed for a linearized system to initialize the controller network parameters, with the controller network selected to be linear (i.e., has no hidden layers) [3]. Inspired by this idea and considering the parameter-dependent dynamics for an NPV system, we warm-start our PD controller by pre-training it to approximate a series of LQR controllers. For each $\theta\in\Theta$, the NPV system in linearized about the trim points $(x^{*},u^{*}(\theta))$. We then minimize the MSE loss $\pi_{\textup{NN}}(x,\theta)-K(\theta)x||^{2}$ over $(x,\theta)$ samples so that this PD controller can somewhat capture the behavior of parameter-dependent LQR controllers.
III-C Empirical Verification
III-C1 PGD-based Verification
III-C2 Trajectory-based Verification
Algorithm 1 Neural-NPV, Parameter-dependent Neural Lyapunov Control
1:Lyapunov network $\phi_{\mathrm{NN}}(\cdot)$; Control network $\pi_{\mathrm{NN}}(\cdot)$; Initial weights & biases for controller network; Learning rate $\eta$; Region of interest $\mathcal{X}\times\Theta$; Counterexample (CEX) buffer $\mathcal{D}$; Candidate sample buffer $\mathcal{C}$; Maximum iteration $N_{\textup{iter}}$; Maximum epoch $N_{\textup{epoch}}$; PGD step $N_{\textup{PGD}}$; Max stagnation $n_{s}$; Samples per iteration $S$; Exponential rate $k_{3}$; PGD step size $\beta$; Loss weights $\beta_{1}$ & $\beta_{2}$; level-set constant $\rho$ (often selected to be 11)
@@ III-C2 Trajectory-based Verification
27:  Sample randomly $S$ points ($x_{\textup{candidate}},\theta_{\textup{candidate}})\in(\mathcal{X}\times\Theta$)
28:  Buffer $\mathcal{C}$ $\leftarrow$ ($x_{\textup{candidate}},\theta_{\textup{candidate}})$
@@
30:   Compute $L_{\dot{V};\rho}$ in 19 using all samples in $\mathcal{D}$
31:   Compute $L_{\Lambda^{\rho}_{V}}$ in 20 using all samples in $\mathcal{C}$
32:   $W_{V},W_{\pi}\leftarrow{W_{V},W_{\pi}}$ $\eta\nabla_{(W_{V},W_{\pi})}\Bigl(\beta_{1}ReLU(L_{\dot{V};\rho})+\beta_{2}ReLU(L_{\Lambda^{\rho}_{V}})\Bigr)$
33:  if CEX violation rate << 1% and no new CEX added to $\mathcal{D}$ for $n_{s}$ iterations then
34:   BREAK
35:Output $\phi_{\textup{NN}}$ & $\pi_{\textup{NN}}$
@@ IV Simulation Results
In this section, we demonstrate the performance and scalability of our proposed methods through simulation experiments using two systems: a simple inverted pendulum and a quadrotor.

### 07 Apr 2026 v2

@@ I Introduction
A nonlinear parameter-varying (NPV) system is a nonlinear system whose behavior explicitly depends on external time-varying parameters. This concept extends the linear parameter-varying (LPV) framework [17], by allowing for the dynamics to have a nonlinear instead of a linear structure. Many real-world nonlinear time-varying systems can be effectively represented as NPV systems. Examples include rockets with decreasing mass due to fuel consumption, ground vehicles subjected to variable friction coefficients, and aircraft experiencing disturbances that change over time. Similar to the LPV framework [17, 12], representing a nonlinear time-varying system as an NPV system facilitates the design of gain-scheduled nonlinear controllers that adjust the control strategy dynamically based on real-time measurements and estimates of the scheduling parameters, while the limitations of the LPV framework resulting from local validity or over-approximation of the LPV models [17].
Recent studies have explored the synthesis of nonlinear parameter-dependent (PD) controllers for NPV systems using SOS programming [8, 10, 26]. However, SOS programming is not scalable to high-dimensional systems. Additionally, SOS-based nonlinear synthesis often imposes a restrictive structure on the Lyapunov functions to avoid a non-convex problem, leading to potentially conservative results. To overcome the limitations of SOS-based NPV synthesis, this paper investigates the use of neural networks to jointly learn PD controllers and Lyapunov functions for NPV systems.
@@ I-A Related work
SOS and its use in NPV synthesis: One of the most widely used methods for nonlinear analysis and synthesis is SOS optimization [13]. It has been applied to jointly synthesize nonlinear parameter-dependent (PD) controllers and PD Lyapunov functions for NPV systems [8, 10, 26]. However, SOS optimization-based methods suffer from two major limitations: (i) Restricted scalability: The computational complexity grows rapidly with the dimensionality of the system and the degree of the polynomial basis used, rendering them impractical for high-dimensional systems [2]. (ii) Conservative performance due to structural restrictions: SOS-based nonlinear synthesis typically imposes structures on certificate functions to avoid a nonconvex problem, leading to conservative performance. For example, SOS-based Lyapunov synthesis typically requires that the Lyapunov functions do not depend on states whose derivatives are directly affected by control inputs [14, 26].
@@
Learning certified control: There has been an increasing trend for learning a control certificate (alongside a controller) using NNs (NNs) [5]. Such a neural certificate can be a Lyapunov function for certifying the stability of an equilibrium point [15, 3], or a contraction metric for certifying stability about trajectories [20, 21] or a barrier function for certifying set invariance [16, 7], among others. Compared to SOS programming-based nonlinear synthesis, certified neural control can be applied to general nonlinear systems (in contrast to polynomial systems for SOS programming-based methods), and is scalable to high-dimensional systems. It also allows for exploring more general forms of controllers and certificates, potentially reducing the conservatism of traditional methods that typically impose special structures on the controllers and/or certificates (e.g., Lyapunov functions) to avoid a non-convex problem [14, 11, 24, 18, 26]. For a comprehensive overview of these methods, readers can refer to the survey paper [5]. Compared to traditional optimization-based methods for control certificate synthesis, such as sum of squares (SOS) programming [13, 14], Nevertheless, existing work typically consider time-invariant dynamics [6]. To the best of our knowledge, there is no study on learning-based control for NPV systems.
@@ II Problem settings and preliminaries
$\theta(t)\in\Theta\;\;\text{and}\;\;\dot{\theta}(t)\in{\Omega},\quad\forall\,t\geq 0,$ (3)
@@
[9, Definition 4.5] The equilibrium point $x^{*}=0$ of the closed-loop system 7 is exponentially stable if there exist positive constants c,kc,~k, and $\lambda$ such that
@@
According to Lyapunov stability theory (as demonstrated in the proof of Theorem 1 in [26]), if there exists a PD controller $u=\pi(x,\theta)$ and a PD continuously differentiable function $V(x,\theta)\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}$ such that for all $x\in\mathcal{X}$ with $\mathcal{X}$ being a region of interest (ROI),
@@
$\dot{V}(x,{\theta},\dot{\theta})\!=\!\frac{\partial V}{\partial x}\cdot\bar{f}+\frac{\partial V}{\partial\theta}\cdot\dot{\theta},$ (10)
@@ III Neural-NPV: Learning PD Controllers and Lyapunov functions
III-A Parameterizing the PD controller and Lyapunov function
@@
$\tilde{u}(x,\theta)=\pi_{\textup{NN}}(x,\theta)-\pi_{\textup{NN}}(x^{*},\theta)+\tilde{u}(x^{*},\theta).$ (14)
. To enforce the constraint 2, we use tanh in the output layer of the controller $\pi_{\textup{NN}}$ to restrict the normalized output within [1,1][-1,1]. $\tilde{u}(x,\theta)$ and $\tilde{u}(x^{*},\theta)$ are normalized value of $u(x,\theta)$ and $u(x^{*},\theta)$, respectively. Rather than using projection or clipping, which introduces complications in gradient-based training, we employ a smooth, differentiable activation at the output, which empirically improves the training process. We scale the normalized output to actual physical control input using an affine de-normalization to get the final controller input:
@@ III-B The Neural-NPV algorithm
Within Neural-NPV, we jointly train the Lyapunov network $\phi_{\textup{NN}}$ and controller network $\pi_{\textup{NN}}$ using a two stage procedure. In Stage I, we use counterexample guided synthesis to obtain a working controller and a Lyapunov function candidate over the full domain $\mathcal{X}\times\Theta$, where $\mathcal{X}$ is the region of interest. In Stage II, we refine the controller and Lyapunov networks using a level-set-guided loss to certify a robust ROA and a surrogate loss that promotes the growth of the robust ROA. The algorithm is summarized in 34.
@@ III-B1 Stage I: Counterexample-guided joint synthesis
Existing work on learning certified control typically first generates samples from the state space, use these samples to evaluate the violation of the Lyapunov functions and optimizes the network parameters to minimize the violation. We follow this idea. However, here, we sample from the space of $x$, and $\theta$, and $\dot{\theta}$ i.e., $\mathcal{X}$, $\Theta$, and $\Omega$. At each iteration, we generate $N$ number of $x$-$\theta$ pairs, and then use CEGIS (Counter-example guided Inductive Synthesis) [1] framework to iteratively generate adversarial samples $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ using projected gradient descent (PGD) [23] to generate the worst adversarial examples that violate the following objective:
$\hskip-11.38109ptL_{\dot{V}}=\sum_{\dot{\theta}\in\textup{Ver}(\Omega)}\!\dot{V}(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}},\dot{\theta})+k_{3}\|x^{i}_{\textup{adv}}\|^{2}.$ (16)
By maximizing 16, we aim to find adversarial samples that have the maximum violation of the condition 9c. Since 9a and 9b has already been satisfied by construction as explained in Section III-A, we only need to evaluate 9c. At each $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$, we evaluate $V(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and $u(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and thus assess whether the conditions in 9 are violated at $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and if so, the level of violation. These adversarial examples are added to a buffer and then used to optimize the following empirical Lyapunov loss:
$\hskip-11.38109ptL_{lya}=\frac{1}{N}\sum_{i=1}^{N}\textup{ReLU}\!\!\!\!\!\!\sum_{\dot{\theta}\in\textup{Ver}(\Omega)}\!\!\!\!\left(\dot{V}(x^{i}_{\textup{adv}},\theta^{i}_{adv},\dot{\theta})+\!k_{3}\|x^{i}_{\textup{adv}}\|^{2}\right),$ (17)
@@ III-B2 Stage II: Level set-guided ROA refinement
$\Lambda^{\rho}_{V}=\left\{x\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\max_{\theta\in\Theta}\,V(x,\theta)\leq\rho\right\},$ (18)
@@
$\hskip-5.69054ptL_{\dot{V};\rho}\!=\frac{1}{N}\sum_{i=1}^{N}\!\min\left(L_{\dot{V}}(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}},\dot{\theta}),\rho\!-\!\max_{\theta\in\Theta}V(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})\right).$ (19)
@@
$\vskip-19.91692ptL_{\Lambda^{\rho}_{V}}=\frac{1}{N}\sum_{i=1}^{N}\max_{\theta\in\Theta}V(x^{i}_{\textup{{candidate}}},\theta^{i}_{\textup{{candidate}}})-\rho.$ (20)
Equation 20 indirectly promotes the growth of the level-set $\Lambda^{\rho}_{V}$. We sample uniformly from the state space to get ($x_{\textup{candidate}},\theta_{\textup{candidate}}$) that are used to compute 20. This surrogate loss ensures the robust level-set $\Lambda^{\rho}_{V}$ does not shrink to a neighborhood of the origin, and cover the region of interest as much as possible. The cost function of this refinement stage consists of both these objective functions. These two objectives play a complementary role, 19 certifies $\Lambda^{\rho}_{V}$ as a valid robust ROA while 20 tries to maximize its volume.
Algorithm 1 Neural-NPV Control
1:Lyapunov network $\phi_{\mathrm{NN}}$; Control network $\pi_{\mathrm{NN}}$ with initial parameters; Learning rate $\eta$; Region of interest $\mathcal{X}\times\Theta$; Counterexample (CEX) buffer $\mathcal{D}$; Candidate sample buffer $\mathcal{C}$; Maximum iteration $N_{\textup{iter}}$; Maximum epoch $N_{\textup{epoch}}$; PGD step $N_{\textup{PGD}}$; Max stagnation $n_{s}$; Samples per iteration $S$; Constant $k_{3}$; PGD step size $\beta$; Loss weights $\beta_{1}$ & $\beta_{2}$; level-set constant $\rho$
@@
27:  Sample randomly $S$ points ($x_{\textup{candidate}},\theta_{\textup{candidate}})\in(\mathcal{X}\times\Theta$), and append to buffer $\mathcal{C}$ $\leftarrow$ ($x_{\textup{candidate}},\theta_{\textup{candidate}})$
@@
29:   Compute $L_{\dot{V};\rho}$ in 19 using all samples in $\mathcal{D}$
30:   Compute $L_{\Lambda^{\rho}_{V}}$ in 20 using all samples in $\mathcal{C}$
31:   $W_{V},W_{\pi}\leftarrow{W_{V},W_{\pi}}$ $\eta\nabla_{(W_{V},W_{\pi})}\Bigl(\beta_{1}\textup{ReLU}(L_{\dot{V};\rho})+\beta_{2}\textup{ReLU}(L_{\Lambda^{\rho}_{V}})\Bigr)$
32:  if CEX violation rate << 1% and no new CEX added to $\mathcal{D}$ for $n_{s}$ iterations then
33:   BREAK
34:Output $\phi_{\textup{NN}}$ & $\pi_{\textup{NN}}$
III-B3 Initialization of the control network
Before entering Stage I, we initialize our PD controller to enhance the training process. This is to avoid convergence issues when starting with a randomly initialized controller network. Inspired by [3], we first design a gain-scheduled linear state-feedback law based on LQR using local linearization at various $\theta$. We then train the initial controller network to match the gain-scheduled controller, i.e., we minimize the loss $\pi_{\textup{NN}}(x,\theta)-K(\theta)x||^{2}$ over $(x,\theta)$ samples so that the PD controller has a reasonable starting point.
III-C Empirical verification
III-C1 PGD-based Verification
III-C2 Trajectory-based Verification
@@ IV Simulation Results
We demonstrate the performance and scalability of our proposed methods through numerical experiments using two systems: a simple inverted pendulum and a quadrotor.